# Planning in a recurrent neural network that plays Sokoban

### v2

@@ 4.1 Probe methodology: training, intervention and targets
We train logistic or linear regression probes with L1 decay to predict various features of the environment from the agent’s activations. We train on two types of inputs. 1) Grid-wise inputs [Bush et al., 2024]: each square in the $10\times 10$ grid is a different data point, potentially with a different label depending on location and timestep. The input is the $64$-dimensional LSTM state $(h,c)$ at a square, usually concatenated for the 3 layers. 2) Global inputs: Each time step produces only one data point. The probes take as input the 64 channels of all of the squares in the 10x10 grid, resulting in a 6400-dimensional input for each layer. In both cases, the per-layer activations are concatenated for probes that are trained on multiple layers.
The dataset consists of states collected by evaluating the $\text{DRC}(3,3)$ on the hard Boxoban levels, excluding the first 5 steps of each episode because the plan is still forming. For multi-class probes, training and F1 scores are computed as one-vs-all: the presence of absence of a particular class is the probe label. We search the best learning rate and L1 decay with grid-search by evaluating the F1 on a validation split of $20\%$ of timesteps from the hard levels. We take the first 1000 medium-validation levels as our test set to report the results (Table 1).
@@
Given a probe that is predictive of future behavior, we edit the agent’s activations such that the probe predicts the behavior we desire to induce. To the extent that the agent carries out the edited behavior, we can conclude that it executes a plan encoded in the linear direction of the probe.
For a linear probe on the activations $h$ with parameter vector $p$, we intervene in order to increase the logits for a particular behavior, given by $p\cdot h$ (we omit a scalar intercept for brevity). Unlike previous work in steering vectors for LLMs [Turner et al., 2023, Rimsky et al., 2023, Li et al., 2024], we found that simply adding a perturbation equal to a fixed multiple of $p$ caused the policy to degenerate and choose random actions. We instead add an adaptively-scaled multiple of $p$ to the activations, chosen to minimize the change in the activations while still causing the probe’s logit to be greater than a desired amount $\alpha$. More precisely, we find $\operatorname*{arg\,min}_{h^{\prime}}\|h^{\prime}-h\|_{2}^{2}$, with the constraint that $h^{\prime}\cdot p>\alpha$. The solution to this is $h^{\prime}=h+{\hat{p}}\max(0,\alpha-(h\cdot p))$, where ${\hat{p}}=p/\|p\|^{2}_{2}$.
Targets.
We choose targets that plausibly encode the steps of a plan. Most of these labels are for grid-wise probes, but the pacing and value probes are global. The direction probes are very similar to Bush et al. [2024].
  • Agent-Directions probe: This probe has 5 outputs: NV (No Visit), UP, DOWN, LEFT, RIGHT. The probe takes as input the 64-dim hidden state activations at a square $(x,y)$ at timestep $t_{i}$. If the agent visits the square $(x,y)$ at a future timestep $t>t_{i}$, then the probe predicts the corresponding direction that the agent takes from that square. For squares that are visited multiple times in the future, the direction corresponding to the first future visit is taken as the target. If a square is never visited in a future timestep, the probe predicts NV. There is one single multinomial probe, for all time steps and locations $(x,y)$.
  • Boxes-Directions probe: This probe is similar to the above probe, except that it is trained to predict the direction in which any of the four boxes will move at a given square in the future. The rationale for this probe is that the boxes, and constraints on their movement given by obstacles, are the main difficulty with Sokoban puzzles.
@@
Probe target Predictive Probe Results (a) Causal Probe Results (b)
Sparse Probes Causal Effect
Best F1 F1 $L_{0}$-norm $\alpha$ Average Best-case
Box-directions $86.4\pm 0.1$ $73.1\pm 0.1$ $63$ 30 $43.7\pm 0.6$ $77.8\pm 0.9$
Agent-directions $72.3\pm 0.1$ $61.3\pm 0.1$ $121$ 10 $7.1\pm 0.3$ $20.7\pm 0.7$
Next box $74.2\pm 0.4$ $69.7\pm 0.5$ $51$ 40 $5.5\pm 1.0$ $15.1\pm 2.5$
Next target $54.3\pm 0.5$ $44.0\pm 0.5$ $32$ 30 $4.6\pm 0.8$ $13.2\pm 2.0$
Pacing $31.0\pm 1.8$ $31.0\pm 1.5$ $5$
@@ 4.2 Probe evaluation and causality
Most probes are quite predictive, as seen in Table 1. The exception is the pacing probe, with $F_{1}=31.0\%$, which is not much better than the constant 1 probe, which has $F_{1}=12.8\%$. This leads us to conclude that the $\text{DRC}(3,3)$ does not represent whether it is in a cycle or not. See Figure 1 (right) and Figure 6 for visualizations of agent-direction and box-direction probes [^5] The supplementary material provides visualization videos of all the probes across several levels..
@@
We measure how causal the probes are by intervening at every step on the hidden state $h,c$ at the layer the probe was trained on, following the procedure in section 4.1 and checking whether the agent follows that action. We report the result in Table 1(b) by performing a grid-search over the strength of intervention $\alpha$ and picking the best value. Despite our adaptive scaling of the intervention, a high value of $\alpha$ can still disturb the agent such that it starts performing random actions whereas a low value of $\alpha$ may not be sufficient to cause the agent to follow the intended behavior. Most of the probes are not causal: only the “boxes directions” probe (used in Figure 1) is causal while the “agent-direction” probe mildly causal.
Even then, we have observed that it can only change boxes directions when they would not deadlock the level in a naive way. That is, if you try to make the model push a box onto a wall, such that the box would not be able to reach a target anymore, it does not. For this reason, we introduce the Best-case causal effect in Table 1(b): try all three directions that are not the actual direction that the box follows, and count the probe as “causal” if it works for any of them (e.g. because they do not naively deadlock the level).
The box-direction probe’s coverage increases during thinking time.
Figure 5 (middle) show that the amount of squares in which the probe predicts a direction, be they lone or linked to a box, grows over thinking steps. Its description of what the box movements will be also greatly improves after a few steps of thinking Figure 5 (right).
4.3 Bistable and unstable plans in toy environments
The causal intervention results from Table 1 show that the box-directions probe is much more causal than the agent-directions probe. We further study this in an isolated setting. Figure 6 (left) shows a level with only a single box next to a target, where the agent has only two equally good paths it can follow. Without intervening, the agent first takes the path on the right by moving three steps, while the agent-direction probe grows towards the box. After the first three moves, the chain connects. The agent then takes a U-turn to follow the path on left. This is very unlike the optimal behavior, which is to commit to any path and follow it.
We can get the agent to commit to one path early by intervening with the agent-direction probe on any of the two paths on the first step only, as shown in Figure 6 (middle). When we construct a similar level where there are two equally good paths for a box to follow, we find that the agent-direction probe is no longer causal; but the boxes direction probe does get the agent to commit. Figure 6 (right) shows an almost empty level. When we intervene with the box-directions probe on the first four steps, the agent follows the laid out path in those steps, but then quickly switches to a more optimal path with fewer turns. The laid out path (not taken, shown in red) is not followed after stopping the intervention.
This analysis provides strong evidence that the network primarily builds upon the boxes directions represented in the activations and relies on the agent-direction when the boxes directions do not inform the subsequent actions.
Figure 6: Left: A custom level where the agent has two equally good paths to follow. The arrows show the prediction of the agent-direction probe with opacity proportional to the number of times an arrow was predicted across all the steps. The green and red arrows are correct and incorrect predictions, respectively. The agent behaves suboptimally by returning to the start and going left after first going right for three steps. Middle: The agent takes the path on the right after intervening with the corresponding arrows using the agent-direction probe on the first step. The same happens on the left if we intervene on that path without the suboptimal steps right of the undisturbed agent. Right: An empty level with box-directions probe intervening on the first four steps which are followed correctly by the agent on those four steps. When the intervention is removed on the fifth step, the agent computes the simpler path in green and doesn’t follow the path laid out earlier (in red).
5 Model surgery exposes latent generalization capabilities
A common feature of search algorithms is that they generalize very well to unusual problems in their domain. This was the main thrust of the argument of Guez et al. [2019]: they exhibited many ways in which the $\text{DRC}(3,3)$ generalized to levels outside of the training distribution to argue that it does search.
5.1 Generalizing beyond $10\times 10$ inputs and training examples
This section goes beyond the capabilities of $\text{DRC}(3,3)$ as a black-box unit. By looking for interpretable features in the activations of the ConvLSTM core, we found that layer 3, at the last tick, has four channels which represent the next action to take. Since the network up until the ConvLSTM core is completely convolutional, we can evaluate it on inputs of any size – unlike $\text{DRC}(3,3)$ as a whole, which only works on $10\times 10$ inputs, and is trained only on levels with four boxes.
@@ 5.1 Generalizing beyond $10\times 10$ inputs and training examples
We start with the Layer 3 grid-wise next-action probes from Table 3. For each spatial location, these predict the action the $\text{DRC}(3,3)$ will take in this step. We attempt three different spatial aggregation methods: mean-pooling, max-pooling, and the proportion of locations with positive probe readouts for a particular action. We learn their relative weights and a baseline for each action by using Adam with learning rate $\ell=10^{-3}$, annealed linearly to 0 for 10000 steps. Table 2 lists the learned parameters, which place most weight on the mean-pooling aggregation rule. The training set consists of 3000 levels, 1000 from each of the Boxoban unfiltered, medium, and hard training sets.
Table 2: Weight and bias for transforming direction probes to predictions
Mean Max Positive proportion
Weight 1.2086 -0.0582 0.2070
Up Down Left Right
Bias 0.3337 -0.0921 -0.0632 -0.0539
@@ Results.
The aggregated features obtain 83% accuracy at predicting the MLP’s action output on the training set, which is less than expected given the $90\%+$ F1 scores of probes in Table 3. With only seven parameters, we did not bother collecting an i.i.d. test set and went straight to the generalization tests.
The adjusted ConvLSTM is able to solve many levels that are out-of-distribution: larger than $10\times 10$ in both dimensions, and with more than four boxes, e.g. those in Figure 2. Out of our test set collected from Þorsteinsson [2009], it can solve $59/480$ levels with both dimensions $>10$, $30/203$ levels with only one dimension $>10$ and $180/482$ levels with both dimensions $\leq 10$. Appendix B has a breakdown by level collection.
5.2 Looking for interpretable features with Sparse Autoencoders
The way we found out actions were represented in the ConvLSTM was with sparse autoencoders (SAEs). We inspected an SAE [Huben et al., 2023, Bricken et al., 2023] trained on the last layer with the top-k activation [Gao et al., 2024, Tamkin et al., 2024], $k=8$. However, we then noticed that the action features were also represented directly in the ConvLSTM channels.
Table 3 compares raw channels, SAE features, and linear probes for how well they match the interpretation of “the feature fires when the $\text{DRC}(3,3)$ takes a particular action”. We find that the raw channels and linear probes are similarly good, and both are considerably better than the SAE.
Thus, for these features, SAEs actually worsened interpretability. We also found three other interpretable features in the SAE, which were also represented directly in the channels (Appendix E).
Feature Channel SAE Feature Probe
Number Prec Rec F1 Number Prec Rec F1 Prec Rec F1
Up L3C29 95.7 88.1 91.7 L3F270 93.9 76.2 84.1 97.5 86.5 91.7
Down L3C8 98.4 80.8 88.8 L3F187 98.0 79.1 87.6 97.6 86.9 91.9
Left L3C27 85.5 84.6 85.1 L3F244 96.1 63.2 76.2 83.5 86.6 85.0
Right L3C3 97.0 86.9 91.7 L3F385 94.6 78.5 85.8 97.6 87.4 92.2

### v3

@@ 3.1 Probe methodology: interpretation and intervention
We train linear regression probes with L1 decay to predict environment features from the agent’s activations. Similar to the concurrent work by Bush et al. [2025], we use two types of probes: (1) Grid-wise inputs treat each square in the $10\times 10$ grid as a different data point, with inputs consisting of the $64$-dimensional LSTM state $(h,c)$ at a square, concatenated for the 3 layers. (2) Global inputs aggregate information from the entire 10x10 grid, resulting in a 6400-dimensional input for each layer at every timestep.
The train and test dataset comprises states collected by evaluating the $\text{DRC}(3,3)$ on the hard Boxoban levels, excluding the first 5 steps of each episode as the plan is still forming. Bush et al. [2025] evaluated their probes on simple toy-levels similar to Appendix D, which don’t require long-term planning for most levels whereas we evaluate our probes on the hard set that can’t be solved greedily and thus require long-term plans. For evaluation of multi-class probes, F1 scores are computed as one-vs-all: the presence or absence of a particular class is the probe label. We search the best learning rate and L1 decay with grid-search by evaluating the F1 on a validation split of $20\%$ of timesteps from the hard levels (Table 1).
Concepts (labels).
We choose concepts that are likely to encode steps in the agent’s plan. Both the Agent-Directions and Box-Directions probes are grid-wise. These are central to our causality analysis and the same as Bush et al. [2025]. We train five additional probes: Next-Box, Next-Target, Next-Action, Pacing, and Value, described in Appendix F.
  • Agent-Directions probe. Predicts the direction the agent takes from a square $(x,y)$ at the nearest future timestep. It has 5 outputs: NV (No Visit), UP, DOWN, LEFT, RIGHT. The probe takes the 192-dim hidden state activations concatenated across the 3 layers at a square $(x,y)$ as input. If the agent visits the square in the future, the probe predicts the corresponding direction; otherwise, it predicts NV. Thus, for each image observation, we get a $10\times 10$ target.
  • Boxes-Directions probe. Same, but predicts the direction any of the four boxes will move in at a given square.
@@
To see if the NN uses these concepts in its algorithm, we edit the network’s activations to influence probe outputs [Li et al., 2023], and measure the causal effect: the fraction of cases where an intervention affects the agent’s action accordingly. [^4] This measurement is known in the literature as the average causal effect or average treatment effect [Holland, 1986], the expected difference in a variable (whether the direction changes) when a causal intervention is present or absent.. For a linear probe on activations $h$ with parameter vector $p$, we intervene to increase the logit $p\cdot h$ to an adaptive value $\alpha$, following work on steering vectors [Turner et al., 2023, Rimsky et al., 2023, Li et al., 2024]. More precisely, we find $\operatorname*{arg\,min}_{h^{\prime}}\|h^{\prime}-h\|_{2}^{2}$, with the constraint that $h^{\prime}\cdot p>\alpha$. The solution to this is $h^{\prime}=h+\hat{p}\max(0,\alpha-(h\cdot p))$, where $\hat{p}=p/\|p\|^{2}_{2}$. For multiclass direction probes, we also zero-out the logits of the other classes. For example, to influence the UP direction at a square, we set the UP logit to $\alpha$ and the DOWN, LEFT, RIGHT, and NV logits to 0. This intervention more precisely edits the network’s representation as compared to the intervention of $h^{\prime}=h+p$ performed by Bush et al. [2025] on toy-levels. Our intervention method is required to observe causal effects on standard difficult levels in the hard set.
@@
Probe Pred (a) Causal Effect (b)
F1 $\alpha$ Avg Best-case
Box-Dir $86.4\pm 0.1$ 30 $49.3\pm 2.0$ $82.5\pm 2.5$
Agent-Dir $72.3\pm 0.1$ 10 $7.1\pm 0.3$ $20.7\pm 0.7$
Next box $74.2\pm 0.4$ 40 $5.5\pm 1.0$ $15.1\pm 2.5$
Next target $54.3\pm 0.5$ 30 $4.6\pm 0.8$ $13.2\pm 2.0$
@@ 3.2 Probe evaluation and causality
Tables 1 and 10 show that most probes demonstrate high predictive accuracy.
Figure 1 (right) and 17 show visualizations of Agent-Directions and Box-Directions probes [^5] The supplementary material provides visualization videos of all the probes across several levels.. In particular, the box-directions probe can predict future box movements up to 50 steps in advance with over 90% accuracy (Figure 4, left).
@@
For each time step, we measure the causal impact of probes by intervening on the hidden state activations $h,c$ at the layer the probe was trained on. The results are reported in Table 1(b), where we perform a grid-search over the intervention strength $\alpha$ to find the optimal value. Despite our adaptive scaling, high values of $\alpha$ can still destabilize the agent’s behavior, causing random actions, while low values of $\alpha$ may fail to induce the intended behavior. Most probes show little causal effect, the Box-Directions probe (used in Figure 1) demonstrates strong causality, and the Agent-Directions probe shows mild causality.
Even then, the Box-Directions probe seems to alter box movements only when they do not lead to naive deadlocks. For instance, the model avoids pushing a box into a wall if later steps indicate that it would need to get the box off the wall to get it to the target. To account for this behavior, we introduce the Best-case causal effect in Table 1(b). This tests all three alternate directions that a box could move, counting the probe as “causal” if it successfully influences any of these directions without causing a naive deadlock.
By intervening on the probes, we can lock the DRC into a plan when there are two equally valued options. Figure 17 in Appendix D shows visualizations of how causal interventions with the directions probe alter the trajectory of the agent. However, in most cases, if we stop intervening with a suboptimal plan, the DRC formulates a better plan online and follows it. This and the naive deadlock problems may be caused by us intervening only on the current square of the box, target or agent: the NN activations still contain contradictory information about other squares, so the NN malfunctions.
4 The plan improves with computation
Guez et al. [2019] showed that the DRC solves more levels if we give it thinking steps at the start of episodes (see Definition 1.4). But what happens in the network during these thinking steps? Using the Box-Directions probe from Section 3, we show that the represented plan increases in length and accuracy (Figure 4, middle and right), demonstrating finding 2. Extra thinking is more helpful for levels that have longer optimal solutions (Figure 3, right), or which require less myopic thinking to place the first few boxes (Figure 3, left)(b). Thus, we are moderately confident that extra thinking improves the success rate by letting the network develop its plan long enough that it finds a non-myopic solution instead of accidentally locking the level.
Figure 3: Left: Average time step to place each box $B_{i}$ on target for different numbers of thinking steps. (a) Averages across all levels. (b) Averages over levels solved by $6$ thinking steps but not solved by 00 thinking steps. More thinking steps make the DRC avoid greedy strategies in favor of long-term return. The 95% confidence intervals computed with the bootstrap method over the levels is very small and not visible in the plot. Right: Average optimal solution length of levels grouped by the number of thinking steps at which the level is first solved. Levels that take longer to solve tend to be harder. NS stands for “not solved”.
Effects of thinking time: non-myopic network
We first examine how additional thinking time impacts the DRC performance by performing thinking steps at the start of an episode, allowing the network to process its hidden state. As shown by Guez et al. [2019], adding 6 thinking steps increases the success rate by $4.7\%$, with a slight drop beyond 16 steps (Figure 7).
This improvement is more pronounced for harder levels, which require longer optimal solutions (Figure 3, right). Notably, thinking time does not correlate with the number of nodes expanded by an A* search, suggesting the DRC works differently (Figure 21).
Thinking reduces early error in levels where placing the first box may block completion. Figure 3 (left) (b) shows that without extra thinking steps, the agent pushes the first box to a target about 4 steps too early, and notably much earlier than average (Figure 3, left). Providing 6 thinking steps prevents this mistake. While some improvement stems from avoiding catastrophic moves, the fact that the average time to push the first box exceeds 6 steps even without forced thinking suggests that the network uses the additional time to form a more complete solution, which changes behavior.
The impact of thinking time varies with training duration and level difficulty. Figure 1 (bottom-left) shows that most of the planning improvement occurs within the first 70M steps. Beyond this point, the planning effect grows a bit more until 200M steps and then stabilizes (hard levels) or slowly decays (medium validation). Based on this, we speculate that the network develops better heuristics for when to think on medium-difficulty puzzles, reducing its default myopia.
Plan improvement.
If DRC refines its plan with extra computation (finding 2), probe-extracted plans should become more predictive of the agent’s actions as more computation time is provided. Figure 4 (right) confirms this effect: the plan initially has low accuracy and rapidly improves as DRC is given more thinking steps. Additionally, Figure 4 (middle) shows that the plans increase in complexity over time, with the length of the plans, defined as sum of length of continuous chains starting from boxes, growing significantly during these additional computation steps.
Figure 4: Left: The number of steps in advance that the Box-Directions probe is able to predict a box move. The probe predicts $\geq 90\%$ of the actions 50 steps before they occur. Middle and Right: The plan quality, as measured by summing length of probe-predicted chain of box-directions starting from boxes and counting squares with positive non-empty predictions, and the F1-score increases over thinking steps (Section 3.1). This suggests that the DRC refines its plan during computation, including across the three ticks per environment step.
5 Purposeful “pacing” to get more computation
On occasion, the DRC exhibits a curious behavior: the agent “paces” in a cycle, returning to the same location, without touching any box. This behavior does not advance the puzzle state, so it is sub-optimal due to the per-step penalty incurred. Why is it still present in the trained network?
We hypothesize (3) that this behavior is incentivized by training, and its purpose is to give the DRC enough time to develop the plan. We show this with multiple lines of evidence, based on observing and intervening on NN behavior, and speed of change in the plan during cycles. We also re-train the $\text{DRC}(3,3)$ with smaller penalties for taking NOOP actions, and show that the total number of cycles stays constant while the number of NOOPs increases. Through all these experiments, we take care to highlight that most cycles are deliberate for plan computation, as opposed to a small fraction of cycles that could be occurring accidentally
Unless otherwise noted, the following experiments are for cycles recorded on medium-validation levels. We merge overlapping cycles, resulting in a total of 13 7021370213\,70213 702 cycles.
Cycles happen early in levels.
In some levels, a single sub-optimal step can render the puzzle unsolvable, so it is very important to develop enough of the plan to prevent that. Section 4 showed that extra thinking steps help in large part by preventing these mishaps, so it makes sense that training would incentivize doing similarly early in a level. Accordingly, Figure 6 (left) shows that most cycles start in the first 5 steps of the episode.
Plans improve much faster in cycles.
If the DRC uses cycles to refine its plan, we expect the plans to be worse at the start of cycles (because they need improvement), and to change faster during the cycle. We check this by looking at the F1 score of box-directions probe at the first step of cycles, and the rate of change in F1 and plan length found with box-directions probe, as measured by number of positive predictions of the probe.
Plan length and quality are highly influenced by whether the agent is close to the beginning or end of a level. To account for this, we pair each step in a cycle with a non-cycle step from a random episode, such that their timesteps $t$ match.
Using the hard-level set, we find that the F1 score at cycle start points is $51.13\%\pm 1.52\%$, compared to $68.13\%\pm 1.45\%$ for non-cycles. Per step, the F1 score improves on average by $0.96\%\pm 0.15\%$ during cycles, versus $0.60\%\pm 0.15\%$ during non-cycles. Finally, the plan length (number of squares with a predicted action) grows by 2.03 squares per step during cycles, compared to 1.37 squares per step during non-cycles.
Figure 23 shows the distribution corresponding to some of these aggregate statistics. While cycle-step improvements definitely skew larger, the distributions has some overlap, which suggests that some cycles could be accidental in which F1 improvement is not larger than non-cycles steps.
Training with smaller NOOP penalty.
The original version of the $\text{DRC}(3,3)$ could not take NOOPs, so it has to pace to get time to think. We train for 800M steps a version of the $\text{DRC}(3,3)$ which gets a smaller penalty for taking a NOOP than for moving. If the DRC is deliberately stalling when the plan is not good enough, we would expect it to use NOOPs when it is stalling deliberately. At the same time, the total number of cycles should stay constant: each step incurs a penalty, so it is still optimal to solve the level as quickly as possible.
Figure 5 shows exactly this effect. The per-move penalty is fixed to $0.1$, and the NOOP penalty varies in $\{0.05,0.07,0.09,0.1\}$. We plot the total number per episode of steps in cycles (including NOOPs) and of NOOPs. As expected, the NOOPs per episode increase with smaller NOOP penalties, and the cycle-steps per episode stay roughly constant (slightly decrease). At the same time, the DRC(3,3) performance stays similar: it goes from solving $70\%$ to $60\%$ of levels. The trend of more cycles getting replaced with NOOPs suggest that most cycles are deliberate, but a small fraction could be accidental that occur anyways. We exclude cycles that happen after the last box is placed in a level because when the DRC sees a level is locked, it goes in cycles until the time limit, and that would make the results above largely depend on the time limit.
Figure 5: Top: Average number of NOOPs and cycle steps per episode, when training with a NOOP action that has less penalty than the movement actions. This makes the network substitute many of the cyclic steps with NOOP steps on average, indicating it is a deliberate behavior. The total number of cycle steps (blue) stays constant. Bottom: percentage of validation-medium levels solved for each network. Error bars show the minimum and the maximum over 3 training seeds.
Cycles can be substituted by NOOPs.
Figure 6 (middle) shows that forcing the model to think for six steps eliminates about 75% of the early cycles. Thus, we can again conclude that most cycles are deliberately performed by the network. We can also directly replace would-be cycles with NOOPs. Just when the DRC would start a cycle of length $N$, we intervene and make it take $N$ thinking steps. Figure 6 (right) shows that these thinking steps largely replace the cycles: in at least $60\%$ of the levels, the DRC followed the same trajectory for a minimum of $30$ steps. For comparison, the median solution length for train-unfiltered is exactly 30 steps.
Figure 6: Left: Histogram of cycle start times on the medium-difficulty validation levels. Middle: Total cycles the agent takes in the first 5 steps across all episodes with $N$ initial thinking steps. Right: Proportion of trajectories that match when replacing $N$-length cycles with $N$ thinking steps, tracked for $x$ steps post-cycle.
6 Generalizing beyond $10\times 10$ inputs and training examples
We extend the DRC’s capabilities by modifying its internal structure to generalize beyond the original $10\times 10$ input size and four-box training distribution. Analysis of the ConvLSTM core reveals that layer 3, at the final tick, encodes four key channels representing the next action provided in Table 10. Since the ConvLSTM is completely convolutional, it can process inputs of arbitrary sizes, unlike the full DRC architecture, which flattens the convolution output and passes through an MLP layer.
@@ 6 Generalizing beyond $10\times 10$ inputs and training examples
Using layer 3 next-action probes (Table 10), we predict at each spatial location the next action of the DRC. To turn the probe’s grid-wise outputs into a single action prediction, we need to aggregate them spatially. We linearly combine three spatial aggregation methods: mean-pooling, max-pooling, and the proportion of positive probe outputs. The optimal combination weights and a bias for each action are learned by minimizing cross-entropy loss against the MLP block using the Adam optimizer on 3000 levels (1000 from each Boxoban training set).
@@ Results.
The aggregated features achieve $83\%$ accuracy on the training set and $77.9\%$ on medium-difficulty validation levels. Although lower than individual probes F1 scores Table 10, these enable the adjusted ConvLSTM to solve many out-of-distribution levels, including larger grids and more than four boxes, e.g., Figure 2 (right) and Figure 15. Specifically, from the test set by Þorsteinsson [2009], it solves $64/484$ levels with both dimensions $>10$, $38/203$ levels with only one dimension $>10$, and $221/486$ levels within the original constraints. A detailed breakdown by level collection of performance and largest level solved is provided in Appendix C. The largest level solved per level-collection is between 3-4x larger in area and number of boxes compared to the $10\times 10$ levels with 4 boxes seen during training.
Our findings show that by aggregating interpretable spatial features, the ConvLSTM generalizes to larger and more complex puzzles without retraining. This suggests that much of the DRC’s planning ability is encoded in the ConvLSTM core, and refining spatial predictions allows for scalable generalization in neural planning systems.
Stress-testing on zig-zag.
How well does this extension to larger levels work? We stress the network by placing it in straightforward (but long) $n\times n$ zig-zag levels (Figure 16). The $\text{DRC}(3,3)$ solves all levels for $n\leq 15$, but fails at $n\geq 16$ no matter how many thinking steps.